From GlueXWiki
Jump to: navigation, search

June 26, 2014 FDC+CDC meeting


  1. Readiness review (Beni, Lubomir)
  2. Gas system [1], [2]
  3. DAQ (Beni)
  4. CDC status (Beni)
  5. Electronics (Chris, Nick)
  6. Engineering (Bill)
    • FDC chiller [3]
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Naomi (CMU), Curtis, Dave, Beni, Chris, Nick, Simon, Bill, Fernando, Eugene, and Lubomir (JLab).

Readiness review

- Curtis prepared (Elton requested) a draft of a commissioning plan for the CDC (linked above). Lubomir and Beni will do the same for the FDC. It doesn't have to be a final version. All the other documents needed for the review are ready.

Gas system

- On Wednesday Scott replaced quickly under pressure, the relief valves on the two mixing tanks (first picture). This morning he replaced also the too regulators outsides, because they had some silicon rubber inside, and installed two filters (for the CO2 and Ar lines) that will remove water and oil contaminations. Fernando was concerned about the silicon. Beni: it works with high rates (i.e. with beam), but we need to start bubbling with alcohol.

- Scott had to stop the gas to the detectors this morning for about 4 hours and for this time the pressure in the CDC dropped almost to the atmospheric (~2-3Pa above). The fourth package dropped to about 39Pa (from 60Pa), all the other FDC packages didn't show significant changes in the pressure. Naomi proposed to look with a borescope camera at the downstream Mylar.

- We started using the Oxygen sensor set-up. There was a small issue: the pressure sensor is supposed to stop the pump if the pressure goes below some limit to prevent under-pressure in the chambers. However, the gas is taken just next to the bubbler and the pressure there varies a lot. To avoid this Beni set the pump to run always, and Dave added in the PLS control to shut down all the valves to the Oxygen sensor if the pressure in any chamber goes below 30Pa.

- We measured the Oxygen on Tuesday and Wednesday using the two sensors and (after some flushing) they gave more or less consistent results: CDC ~3-5ppm, FDC1 and FDC4 ~5-8ppm (extremely low values!). Bill is concerned that the sensors may need calibration. Beni: let's see first if we have problems (charge not constant with the drift time). Today Lubomir did measurements right after the gas was restored: CDC ~8ppm, FDC1 ~35ppm, FDC2 ~120ppm, FDC3 ~46ppm, and FDC4 ~25ppm. Very strangely, the chambers that had bigger leakage this morning showed smaller Oxygen contamination.

CDC noise

Long discussions about the noise (first picture above) on the CDC that Beni studied:

- higher noise on higher channel number (Naomi observed the same effect)

- the noise completely disappears only if the HVB card is disconnected from the detector (last picture)

- the noise reduces gradually by turning off more HV/LV channels but still remains even with the FDC HV also off (second picture)

- no change in the noise with additional HVB shielding or by changing the return/ground connections on the HV modules

More studies needed; Fernando will talk to Cody to use the pulser on the fADC125.


- The mesh needs to be connected to ground because of safety reasons: may charge up. Bill will look for opportunities to do this.

- Fernando had discussions with CAEN about the 1550P issues. They are concerned, as well, and will not charge us for the repair (even modules not under warranty). Most of the problems due to bad QA/QC on their side. On our side, we need better characterization of the problem before shipping the module for repair. Example: one of the module had hardware threshold set to very low value, that's why it didn't rump up. Some found this strange, but has to be checked. Beni: one of the modules that was repaired and worked a week ago, showed a problem again, can't rump up!

- Dave needs Chris to do four more boxes (as for FCAL) for temperature and humidity measurements. Fernando prefers to have also some kind of flowmeter on the cooling duct, in addition to the thermocouple.


- Bill and Casey opened the chiller (picture attached). The change of the level we see is on the sub-tank. The chiller is filled through the sub-tank. The fact that the weight didn't change over a month means Fluorinert moves from the sub-tank to the circulating tank that looked almost full. Bill tested the two float switches and one of them didn't work: if that's the case it will not shut off the pump when there's no coolant. Bill will discuss this with the company and probably will do more checks. In the mean time we continue to use the chiller, as it is, and will check the weight again in a month.